OpenStudy (anonymous):

After 252 days, a 12.0-g starting sample of the radioisotope scandium-46 contains only 1.5 g of the isotope. What is the half-life of scandium-46? Explain your reasoning.

OpenStudy (abb0t):

You can pretty much figure this out on your own using the half-life formula: \(\sf \large\color{red}{A=A_0(\frac{1}{2})^{\frac{t}{h}}}\) Where: t = time h = half-life A\(_0\) = initial amount A = final amount the \(\sf \frac{1}{2}\) is the split factor. After a half-life, one pound becomes half a pound..and so forth.